If 6 parts of mango juice are mixed with 10 parts of water, what part of the mixture is mango juice?
I answered 3/5. Is it right?

Answer by ramkikk66(644)


3/5 is not right.

Because the mixture actually has 16 parts, of which 6 parts are juice and 10 parts
are water.

So the part of the mixture that is mango juice = 6/16 = 3/8 = 0.375

Hope you got it now :)